A sports researcher is interested in determining if there is a relationship between the number of home team and visiting team wins and different sports. A random sample of 526 games is selected and the results are given below. Find the expected frequency for E 2,2 to test the claim that the number of home team and visiting team wins are independent of the sport. Round to the nearest tenth if necessary.

Football Basketball Soccer Baseball Total
Home Team 38 153 27 85
Visiting Team 33 96 18 77
Total
Answer #1
Expected frequency of a cell = sum of row*sum of column / total sum
Expected Frequencies
Football Basketball Soccer Baseball Total
Home Team 71*303/527=40.8 249*303/527=143.2 45*303/527=25.9 162*303/527=93.1 303
Visiting Team 71*224/527=30.18 249*224/527=105.8 45*224/527=19.1 162*224/527=68.9 224

Please revert in case of any doubt.

Please upvote. Thanks in advance